OpenStudy (anonymous):

A fox can run at a speed of 42 miles per hour. How many feet per second is this? a. 1,232 ft/s b. 61.6 ft/s c. 3,696 ft/s d. 42 ft/s

OpenStudy (jdoe0001):

how many feet are there in 1 mile?

OpenStudy (candy13106):

um i dont think i know

OpenStudy (candy13106):

its me joshepe its my old account cuz that other one was to slow

OpenStudy (candy13106):

u there

OpenStudy (lucaz):

5280ft=1mile

OpenStudy (candy13106):

ok ok than wat

OpenStudy (jdoe0001):

ohh ahe... yeah... the site is getting laggy anyhow as @lucaz said, 5280 feet in a mile how many seconds in 1 hr?

OpenStudy (candy13106):

60 secondes in 1 min

OpenStudy (lucaz):

60*60s=1h

OpenStudy (candy13106):

yea i know that

OpenStudy (lucaz):

so you have 42(5,280)/60*60 ft/s

OpenStudy (lucaz):

I guess x_X

OpenStudy (candy13106):

well i ned help now plz someone plz

OpenStudy (jdoe0001):

ok... then \(\bf \cfrac{42\cancel{mi}}{\cancel{hr}} \times \cfrac{5280ft}{\cancel{mi}} \times \cfrac{\cancel{hr}}{3600sec}\)

OpenStudy (candy13106):

ok than wats next

OpenStudy (jdoe0001):

well, what did that give you?

OpenStudy (candy13106):

um well ur telling me to find out the answer

OpenStudy (jdoe0001):

well... we're assuming you're meant to get an answer for the exercise the exercise won't be meaningful if the person didn't exercise

OpenStudy (candy13106):

okwell i right this right

OpenStudy (candy13106):